# Verina: Benchmarking Verifiable Code Generation

### v1

@@ 3.2 Benchmark Construction and Quality Assurance
Verina consists of two subsets sourced from different origins: Verina-basic and Verina-adv. For Verina-basic, the maximum (median) LoC for code and specification are 26 (6) and 17 (2), respectively. For Verina-adv, they are 38 (16) and 62 (7). This underscores the diversity and varying difficulty levels between Verina-basic and Verina-adv. We employ a meticulous data curation process that combines careful translation, thorough manual review, and automated mechanisms, leading to a rigorous and high-quality benchmark for verifiable code generation.
Verina-basic: translation from human-written Dafny code. We first consider MBPP-DFY-50 [18], which contains MBPP [24] coding problems paired with human-verified solutions in Dafny. Each instance contains a natural language problem description, code implementation, specifications, proof, and test cases. We manually translated 49 problems into Lean, refining and verifying each translation. To extend the benchmark, we added 59 more human-authored Dafny instances from CloverBench [12]. These were translated into Lean using OpenAI o3-mini with few-shot prompting based on our manual translations, followed by manual inspection and correction. Overall, the coding difficulty in Verina-basic, abstracting away language differences, is comparable to MBPP.
Verina-adv: writing Lean code from scratch. Verina-adv enhances the diversity of Verina by incorporating more advanced coding problems and solutions. They were adapted from student submissions to a lab assignment in a course on theorem proving and program verification. Students, both undergraduate and graduate, were encouraged to source problems from platforms like LeetCode or more challenging datasets such as LiveCodeBench [1]. They formalized and solved these problems in Lean, providing all necessary elements in Verina’s format (Section 3.1). We carefully selected the most suitable and high-quality submissions, resulting in 81 benchmark instances. In addition, we manually reviewed and edited the submissions to ensure their correctness.
@@
  • Necessary negative tests: We mutated each positive test case to construct at least three different negative tests that violate either the pre- or the post-condition, except when the function’s output has boolean type, in which case only a single negative test can be created. We made sure that our ground truth code and specifications do not pass these negative tests.
@@ 4.1 Foundational Tasks and Metrics
Specification generation (SpecGen). Given a description, signature, and optionally code implementation, the model generates a formal specification. Specifications must accurately capture human intent. Let $\phi$ denote the set of correct programs that satisfy human intent and $\hat{\phi}$ the set that aligns with the generated specification. An ideal specification should achieve $\hat{\phi}=\phi$, which entails two properties—(i) soundness ($\hat{\phi}\subseteq\phi$): it is “small enough” to cover only correct programs, and (ii) completeness ($\phi\subseteq\hat{\phi}$): it is “large enough” to cover all correct programs.
In practice, two challenges arise for evaluating $\hat{\phi}$. First, we must capture $\phi$ formally. Verina addresses this by leveraging high-quality ground truth specifications (see Section 3.2) and comprehensive test suites. Second, we need to assess the relationship between $\hat{\phi}$ and $\phi$ to establish soundness and completeness. Since specifications consist of pre-conditions and post-conditions, let $P$ and $\hat{P}$ denote the ground truth and model-generated pre-conditions, respectively, and $Q$ and $\hat{Q}$ the corresponding post-conditions. In Verina, we define the soundness and completeness of $\hat{P}$ and $\hat{Q}$ as follows:
@@
To evaluate SpecGen, we need automatic and robust mechanisms to check if the above relationships hold. Formally proving them is difficult, as they may contain nested quantifiers and complex program properties. LLM-based provers are ineffective in the verification domain, as shown in Section 5, making them unreliable for this use case. Another approach is to convert these relationships into ATP; however, existing tools do not adequately model the necessary Lean features [43]. To overcome these limitations, we leverage a practical testing-based evaluation framework using our comprehensive test suites, as shown in Figure 3. We formalize a given soundness or completeness relationship, denoted by $R$, in Lean. Instead of proving $R$ for universally quantified input and output variables, we check $R$ against concrete values in test cases. For example, to evaluate $\hat{Q}$’s soundness, we check if $P{}(\overline{x}{}){}\,\land\,\hat{Q}{}(\overline{x}{},y{}){}\Rightarrow Q{}(\overline{x}{},y{}){}$ holds for all test cases $(\overline{x}{},y{})$ in our test suite. We denote this simplified version of $R$ as $R^{\prime}$. For many cases, e.g., the specification in Figure 1, Lean can automatically determine if $R^{\prime}$ holds [44] and we return the corresponding result. Otherwise, we employ property-based testing with the plausible tactic in Lean [45]. It generates diverse inputs specifically targeting the remaining universally and existentially quantified variables in $R^{\prime}$, systematically exploring the space of possible values to test $R^{\prime}$. In Section A.4, we provide a detailed description on how we implement these metrics in Lean.
@@
Since our evaluator is based on testing, it can prove that $R$ does not hold through counterexamples, as highlighted in green in Figure 3. While it cannot formally establish $R$ holds, it remains highly robust in this regard, due to our comprehensive test suite with both positive and negative tests, which achieve full coverage on ground truth code implementations. Lean’s property-based testing cannot handle a small number of complicated relationships, for which our evaluator returns unknown. To further enhance the accuracy of our metric, we repeat our evaluation framework in Figure 3 to check $\neg R{}$. We compare the evaluator outcomes on $R$ and $\neg R{}$, and select the more accurate result as the final output. Our final metrics for SpecGen include individual pass@$k$ [3] scores for soundness and completeness of all generated pre-conditions and post-conditions, as well as aggregated scores that soundness and completeness hold simultaneously for pre-condition, post-condition, and the complete specification. Since the evaluation of the specification may return unknown, we plot error bars indicating the lower bound (treating unknown as $R$ does not hold) and upper bound (treating as $R$ holds).
@@
To illustrate our metric, consider the ground truth pre-condition k < s.size at Line 12 of Figure 1, and model-generated pre-condition k < s.size - 1 and k < s.size + 1. k < s.size - 1 can be determined as unsound using the positive test (s : #[1, 2, 3, 4, 5]) (k : 4), while k < s.size + 1 is incomplete based on the negative test (s : #[1, 2, 3, 4, 5]) (k : 5). We more examples of our metrics for specification generation in Appendix C.
@@
Code generation (CodeGen). Given a natural language description, function signature, and optionally specification, the model generates code implementing the desired functionality. Following standard practice, we evaluate the generated code by running it against positive test cases in Verina and reporting the pass@$k$ metric defined by Chen et al. [3]. In Section 4.2, we will explore evaluating the code by proving its correctness with respect to the formal specification.
@@ 4.2 Task Combinations
  • Specification-Guided Code Generation: Given a natural language description, function signature, and the ground truth specification, the model first generates the code and then proves that the code satisfies the specification. This aligns with tasks explored in FVAPPS [17] and AlphaVerus [15].
  • Specification Inference from Code: In some cases, developers may have the code implementation and want the model to annotate it with a formal specification and prove their alignment. This corresponds to the setting in AutoSpec [27], SpecGen [28], and SAFE [16].
  • End-to-End Verifiable Code Generation: For an even higher degree of automation, developers might start with only a high-level problem description in natural language and instruct the model to generate code and specification independently, and then generate the proof. This captures the scenario in Dafny-Synthesis [18] and Clover [12].
In these task combinations, a crucial design consideration is the dependency between code and specification. For example, in specification-guided code generation, it is important to assess how beneficial the ground truth specification is beyond the natural language description, which already captures the developer’s intent. Additionally, for end-to-end verifiable code generation, it is essential to decide the order of the CodeGen and SpecGen modules—whether to make SpecGen dependent on the output of CodeGen, place SpecGen before CodeGen, or run them independently (as in Figure 4). We experimentally explore these design choices using Verina in Section 5.
In end-to-end verifiable code generation, it is crucial that the model generates code and specification independently, rather than sequentially. Otherwise, it may exploit shortcuts by producing definitionally equivalent code-specification pairs, making the proof task trivial. When designing Verina’s tasks, we enforce independence—for example, the model cannot access the code when generating the specification. While we do not yet check for definitional equivalence (e.g., using BEq [46]), we leave this as an important direction for future work.
@@ 5 Experimental Evaluation
Experimental setup. We evaluate a diverse set of nine state-of-the-art LLMs on Verina. We leverage 2-shot prompting to enhance output format adherence, with the 2-shot examples excluded from the final benchmark. For each task, we primarily report the pass@$1$ metric [3]. We provide detailed input prompts, output formats, and LLM setups in Appendix A.
@@
Figure 5: pass@$1$ performance of LLMs on Verina’s three foundational tasks.
@@
All foundational tasks are challenging, especially ProofGen. Figure 5 shows a clear difficulty hierarchy across the three foundational tasks. Code generation achieves the highest success rates across models, followed by specification generation, while proof generation remains the most challenging with pass@$1$ rates below 3.6% for all models. All three tasks pose significant challenges for current LLMs, with constructing Lean proofs that the implementation satisfies the specification being particularly hard and requiring specialized theorem proving capabilities. This also means that for any combined task involving ProofGen, LLMs’ performance will be heavily bottlenecked by the ProofGen subtask. Among the evaluated models, o4-mini, GPT 4.1, Claude Sonnet 3.7, and Gemini 2.5 Flash demonstrate relatively stronger performance across tasks. We report detailed results on pre-condition and post-condition soundness and completeness in Appendix B, where we observe that generating sound and complete post-conditions is generally more difficult than pre-conditions.
Figure 6: pass@$1$ performance on three foundational tasks for Verina-basic and Verina-adv.
Verina-adv is much more challenging than Verina-basic. The comparison between Verina-basic and Verina-adv in Figure 6 reveals substantial difficulty gaps on all three tasks. This demonstrates that problem complexity significantly impacts all aspects of verifiable code generation, and Verina-adv provides a valuable challenge for advancing future research in this domain.
@@
Figure 7: pass@$k$ performance of selective LLMs on the ProofGen tasks in Verina using proof refinement (first row) and direct generation (second row).
@@
Iterative proof refinement shows meaningful improvements. For ProofGen task, besides pass@$1$, we also extend the evaluation of the 4 best performing LLMs (o4-mini, GPT 4.1, Claude Sonnet 3.7, Gemini 2.5 Flash) to further investigate LLMs’ theorem proving capabilities. We evaluate them with iterative proof refinement, where the evaluated model receives Lean verifier error messages and is prompted to revise its proof, and with direct generation, where the evaluated model generates responses independently in each iteration. For both method, we report pass@$k$, the success rate after $k$ rounds of iterations, for $k$ up 64. This metric investigates how much additional interaction helps repair the proof that a single-pass generation would miss, and whether providing Lean verifier feedback improves success rates compared to independent generation attempts.
As shown in Figure 7, iterative proof refinement yields meaningful improvements on simpler problems, with o4-mini improving from 7.41% to 22.22% on Verina-basic after 64 iterations. However, these gains are substantially less on Verina-adv (1.23% to 6.17%), indicating that naive refinement strategies are insufficient for complex proving tasks. Furthermore, comparing refinements and direct generation without error messages demonstrates the clear value of Lean verifier feedback.
Figure 8: Impact of contextual information (reference code or specification input) on CodeGen and SpecGen performance.
@@
Providing ground truth specification benefits CodeGen. Providing ground truth specifications as context consistently improves CodeGen performance across models. Since the ground truth specifications cannot be used directly as code (as explained in  3.2), all CodeGen improvements rely on semantic understanding of the reference specification. On the contrary, providing ground truth code as context shows minimal or negative improvement for SpecGen. While it is possible for LLMs to directly use the ground truth code in the specification, manual inspection of our evaluation results reveals no evidence of such behaviors. This is likely because using code as specification is uncommon in standard development practices, and our prompts A.3 ask LLMs to focus on constraining code behavior rather than replicating implementation details. The asymmetry in using ground truth information for CodeGen versus SpecGen suggests that formal specifications effectively constrain and guide code synthesis, while verbose code implementations may introduce noise to or over-constrain specification generation rather than providing helpful guidance. Furthermore, when using LLM-generated code or specifications as context instead of ground truth, performance generally degrades. The generated artifacts can be of insufficient quality to serve as reliable reference. This suggests that combined tasks, where LLMs must generate both code and specifications jointly, might be significantly more challenging than individual tasks in isolation.
@@ 6 Conclusion and Discussion
Limitations and future work. Despite advancing the state-of-the-art in benchmarking verifiable code generation, Verina has several limitations. First, its size (189 examples) is modest, scaling to a larger dataset suitable for finetuning likely requires automated annotation with LLM assistance. Second, it emphasizes simple, standalone tasks—well-suited for benchmarking but not fully representative of complex real-world verification projects [47, 10]. Third, while existing provers have limited capabilities for handling complex soundness/completeness relationships, our SpecGen metric (Section 4.1) could be improved by potential more capable provers in the future, including those based on LLMs or SMT solvers, to prove soundness/completeness relationships. Finally, while Lean programs in Verina are newly written, the underlying task topics are drawn from widely used sources, posing a risk of data contamination.

### v2

@@ 3.2 Benchmark Construction and Quality Assurance
Verina consists of 189 problems sourced from different origins. We employ a meticulous data curation process that combines careful translation, thorough manual review, and automated mechanisms, leading to a rigorous and high-quality benchmark for verifiable code generation.
To construct Verina, we first consider MBPP-DFY-50 (Misu et al., 2024) as our data source. It consists of MBPP (Austin et al., 2021) coding problems paired with human-verified solutions in Dafny. Each instance contains a natural language problem description, code implementation, specifications, proof, and test cases. We manually translated 49 problems into Lean, refining and verifying each translation. To extend the benchmark, we added 59 more human-authored Dafny instances from CloverBench (Sun et al., 2024). These were translated into Lean using OpenAI o3-mini with few-shot prompting based on our manual translations, followed by manual inspection and correction.
Additionally, Verina incorporates problems adapted from student submissions to a lab assignment in a course on theorem proving and program verification. Students, both undergraduate and graduate, were encouraged to source problems from platforms like LeetCode or more challenging datasets such as LiveCodeBench (Jain et al., 2025). They formalized and solved these problems in Lean, providing all necessary elements in Verina’s format (Section 3.1). We carefully selected the most suitable and high-quality submissions, resulting in 81 benchmark instances. In addition, we manually reviewed and edited the submissions to ensure their correctness.
During our evaluation, we observe problems adapted from student submissions are generally more difficult than problems translated from Dafny datasets on all models, with detailed analysis provided in Appendix B.
@@
  • Necessary negative tests: We mutated each positive test case to construct at least three different negative tests that violate either the pre‐ or the post‐condition, except when the function’s output has boolean type, in which case only a single negative test can be created. We made sure that our ground truth code and specifications do not pass these negative tests.
@@ 4.1 Foundational Tasks and Metrics
Specification generation (SpecGen). Given a description, signature, and optionally code implementation, the model generates a formal specification. Next, we formally define the soundness and completeness relationships between the generated specification and the ground truth specification. Then, we describe our multi-stage evaluation pipeline to assess whether these relationships hold.
Let $\phi$ denote the set of programs that satisfy the ground truth specification and $\hat{\phi}$ the set that align with the generated specification. An ideal generated specification should achieve $\hat{\phi}=\phi$, which entails two properties—(i) soundness ($\hat{\phi}\subseteq\phi$): it is “small enough” to cover only correct programs, and (ii) completeness ($\phi\subseteq\hat{\phi}$): it is “large enough” to cover all correct programs. Since specifications consist of pre-conditions and post-conditions, let $P$ and $\hat{P}$ denote the ground truth and model-generated pre-conditions, respectively, and $Q$ and $\hat{Q}$ the corresponding post-conditions. In Verina, we define the soundness and completeness of $\hat{P}$ and $\hat{Q}$ as follows:
@@
To practically and reliably assess whether the above relationships hold, we develop a multi-stage evaluator based on theorem proving and comprehensive testing, as shown in Figure 3. We denote a given soundness or completeness relationship by $R$. The evaluator first attempts to prove $R$ using LLM-based theorem provers. When the prover is inconclusive, e.g. due to complex quantifier structures or incapability of current LLM-based provers (as detailed in Section A.5), the evaluator proceeds with a practical testing-based framework using our comprehensive test suites. In this testing-based process, we check $R$ against concrete values in test cases.
@@
For example, to evaluate $\hat{Q}$’s soundness, we check if $P{}(\overline{x}{}){}\,\land\,\hat{Q}{}(\overline{x}{},y{}){}\Rightarrow Q{}(\overline{x}{},y{}){}$ holds for all test cases $(\overline{x}{},y{})$ in our test suite. We denote this simplified version of $R$ as $R^{\prime}$. For many cases, e.g., the specification in Figure 1, Lean can automatically determine if $R^{\prime}$ holds (Selsam et al., 2020) and we return the corresponding result. Otherwise, we employ property-based testing with the plausible tactic in Lean (Lean Prover Community, 2024). It generates diverse inputs specifically targeting the remaining universally and existentially quantified variables in $R^{\prime}$, systematically exploring the space of possible values to test $R^{\prime}$. In Section A.5, we provide a detailed description of how we implement these metrics in Lean.
@@
Since our evaluator integrates proof and testing, it can certify $R$ holds when a formal proof of $R$ succeeds, and it can certify $R$ does not hold by producing counterexamples. When only testing passes without a proof, the evaluator returns $R$ might hold, reflecting strong empirical evidence that $R$ holds. While it cannot formally establish $R$ holds, it remains highly robust in this regard, due to our comprehensive test suite with both positive and negative tests, which achieve full coverage on ground truth code implementations. Lean’s property-based testing cannot handle a small number of complicated relationships on some testcases, for which our evaluator returns unknown. To further enhance the accuracy of our metric, we repeat our evaluation framework in Figure 3 to check $\neg R{}$. We compare the evaluator outcomes on $R$ and $\neg R{}$, and select the more accurate result as the final output.
Our final metrics for SpecGen include individual pass@$k$ scores (Chen et al., 2021) for soundness and completeness of all generated pre-conditions and post-conditions, as well as aggregated scores that soundness and completeness hold simultaneously for pre-condition, post-condition, and the complete specification. Since our specification evalutor may return unknown, we plot error bars indicating the lower bound (treating unknown as $R$ does not hold) and upper bound (treating as $R$ holds).
@@
To illustrate our metric, consider the ground truth pre-condition k < s.size at Line 12 of Figure 1, and model-generated pre-condition k < s.size - 1 and k < s.size + 1. k < s.size - 1 can be determined as unsound using the positive test (s : #[1, 2, 3, 4, 5]) (k : 4), while k < s.size + 1 is incomplete based on the negative test (s : #[1, 2, 3, 4, 5]) (k : 5). We provide more examples of our metrics for specification generation in Appendix C.
@@
Code generation (CodeGen). Given a natural language description, function signature, and optionally specification, the model generates code implementing the desired functionality. Following standard practice, we evaluate the generated code by running it against positive test cases in Verina and reporting the pass@$k$ metric defined by Chen et al. (2021). In Section 4.2, we will explore evaluating the code by proving its correctness with respect to the formal specification.
@@ 4.2 Task Combinations
  • Specification-Guided Code Generation: Given a natural language description, function signature, and the ground truth specification, the model first generates the code and then proves that the code satisfies the specification. This aligns with tasks explored in FVAPPS (Dougherty & Mehta, 2025) and AlphaVerus (Aggarwal et al., 2024).
  • Specification Inference from Code: Developers may have the code implementation and want the model to annotate it with a formal specification and prove their alignment. This corresponds to the setting in AutoSpec (Wen et al., 2024), SpecGen (Ma et al., 2025), and SAFE (Chen et al., 2024).
  • End-to-End Verifiable Code Generation: For an even higher degree of automation, developers might start with only a high-level problem description in natural language and instruct the model to generate code and specification independently, and then generate the proof. This captures the scenario in Dafny-Synthesis (Misu et al., 2024) and Clover (Sun et al., 2024).
In these task combinations, a crucial design consideration is the dependency between code and specification. For example, in specification-guided code generation, it is important to assess how beneficial the ground truth specification is beyond the natural language description, which already captures the developer’s intent. Additionally, for end-to-end verifiable code generation, it is essential to decide the order of the CodeGen and SpecGen modules—whether to make SpecGen dependent on the output of CodeGen, place SpecGen before CodeGen, or run them independently (as in Figure 4). We experimentally explore these design choices using Verina in Section 5. Concurrent with our work, CLEVER (Thakur et al., 2025) introduces 161 manually crafted problems sourced from HumanEval (Chen et al., 2021) with ground truth specifications. However, CLEVER only supports the SpecGen task and the specification-guided code generation setting and cannot capture the full spectrum of workflows that Verina enables through both individual and compositional tasks. We provide detailed comparison in Section A.4.
@@ 5 Experimental Evaluation
Experimental setup. We evaluate a diverse set of eight state-of-the-art general-purpose LLMs and three LLMs or agentic frameworks specialized in theorem proving. We leverage 2-shot prompting to enhance output format adherence, with the 2-shot examples excluded from the final benchmark. For each task, we primarily report the pass@$1$ metric (Chen et al., 2021). We provide detailed input prompts, output formats, and LLM setups in Appendix A.
@@
Figure 5: pass@11 performance of LLMs on Verina’s three foundational tasks.
@@
All foundational tasks are challenging, especially ProofGen. As shown in Figure 5, code generation generally achieves the highest success rates across models, followed by specification generation, while proof generation remains the most challenging with pass@$1$ rates below 3.6% for all models. All three tasks pose significant challenges for current general purpose LLMs, with constructing Lean proofs that the implementation satisfies the specification being particularly hard and requiring specialized theorem proving capabilities. This also means that for any combined task involving ProofGen, e.g., the ones in Section 4.2, LLMs’ performance will be heavily bottlenecked by the ProofGen subtask. Among the evaluated models, o4-mini, GPT 4.1, Claude Sonnet 3.7, and Gemini 2.5 Flash demonstrate relatively stronger performance across tasks. We report detailed results on pre-condition and post-condition soundness and completeness in Appendix B, where we observe that generating sound and complete post-conditions is generally more difficult than pre-conditions.
Figure 6: pass@11 for ProofGen across models and proving agent.
@@
Figure 7: pass@$k$ performance of selective LLMs on ProofGen using proof refinement (left) and direct generation (right).
Specialized provers and agentic methods improve proof success rate.
Given the limitations of general-purpose LLMs, we extend our evaluation to specialized theorem-proving models and agentic approaches. As shown in Figure 7, Goedel Prover V2 32B (Lin et al., 2025) and DeepSeek Prover V2 7B (Ren et al., 2025) achieve higher proof success rates compared to general-purpose models. We further evaluate Copra (Thakur et al., 2023), an agentic theorem-proving framework based on tree-search. We use o4-mini as the backbone model and allow at most 64 LLM queries for each sample. Copra demonstrates clear improvements over direct single-pass generation.
@@
Iterative proof refinement shows meaningful improvements. For ProofGen task, besides pass@$1$, we also extend the evaluation of the four strongest general-purpose models (o4-mini, GPT 4.1, Claude Sonnet 3.7, Gemini 2.5 Flash) alongside two specialized LLM-provers (Goedel Prover V2 32B (Lin et al., 2025) and DeepSeek Prover V2 7B (Ren et al., 2025)). We evaluate them with iterative proof refinement, where the evaluated model receives Lean verifier error messages and is prompted to revise its proof, and with direct generation, where the evaluated model generates responses independently without Lean feedback in each iteration. For all methods, we report pass@$k$, the success rate after $k$ rounds of iterations, for $k$ up 64. This metric investigates how much additional interaction helps repair the proof that a single-pass generation would miss, and whether providing Lean verifier feedback improves success rates compared to independent generation attempts.
As shown in Figure 7, iterative proof refinement reliably outperforms direct generation at matched query budgets on both general purpose and proof-specific models, underscoring the value of Lean verifier feedback. A detailed breakdown by problem difficulty is provided in Appendix B.
Figure 8: Impact of contextual information on CodeGen and SpecGen performance.
@@
Providing ground truth specification benefits CodeGen. Providing ground truth specifications as context consistently improves CodeGen performance across models. Since the ground truth specifications cannot be used directly as code (as explained in  3.2), all CodeGen improvements rely on semantic understanding of the reference specification. On the contrary, providing ground truth code as context shows minimal or negative improvement for SpecGen. While it is possible for LLMs to directly use the ground truth code in the specification, manual inspection of our evaluation results reveals no evidence of such behaviors. This is likely because using code as specification is uncommon in standard development practices, and our prompts A.3 ask LLMs to focus on constraining code behavior rather than replicating implementation details. The asymmetry in using ground truth information for CodeGen versus SpecGen suggests that formal specifications effectively constrain and guide code synthesis, while verbose code implementations may introduce noise to or over-constrain specification generation rather than providing helpful guidance. Moreover, replacing ground truth with LLM-generated artifacts generally degrades performance, indicating that combined tasks are more challenging than individual tasks.
@@ 6 Conclusion and Discussion
Limitations and future work. Despite advancing the state-of-the-art in benchmarking verifiable code generation, Verina has several limitations. First, its size (189 examples) is modest, scaling to a larger dataset suitable for finetuning likely requires automated annotation with LLM assistance. Second, it emphasizes simple, standalone coding problems, which is well-suited for benchmarking but not fully representative of complex real-world verification projects (Klein et al., 2009; Leroy et al., 2016). Third, while our current evaluation pipeline overcomes the limitation of current LLM theorem provers using comprehensive testing, the future advances in LLM theorem prover capabilities can enable stronger formal guarantees. Finally, while Lean programs in Verina are newly written, the underlying task topics are drawn from widely used sources, posing a risk of data contamination.
Ethics Statement
We adhere to the ICLR Code of Ethics and ensure compliance with all relevant dataset licenses, as detailed in Section A.1. All data used in this work are publicly available and collected strictly for academic research purposes with proper citation and attribution.
Reproducibility Statement
We are committed to ensuring the reproducibility of our work. All code, benchmark datasets, and evaluation pipelines introduced in this paper are included in the supplementary materials, accompanied by detailed instructions for setup and usage. The dataset construction processes are described in Section 3.2. The evaluation metrics are described in Section 4. Additional implementation details and experimental settings are described in the appendix.